OpenStudy (amanda0618):

WILL GIVE MEDAL The rectangle has an area of 56 m^2. What is the value of x?

sammixboo (sammixboo):

Anymore info?

OpenStudy (amanda0618):



OpenStudy (amanda0618):

heres the rectangle @sammixboo

OpenStudy (amanda0618):

PLEASE HELP SOMEBODY

OpenStudy (ahsome):

Yay, cube x...

rishavraj (rishavraj):

area of rectangle = l*b so \[(x) \times (3x^2 + 2) =56\]

OpenStudy (ahsome):

\[x(3x^2+2)=56\]\[3x^3+2x=56\]

OpenStudy (amanda0618):

whats next after that? @rishavraj

rishavraj (rishavraj):

just solve the equation...it would be \[3x^3 + 2x = 56\]

OpenStudy (amanda0618):

but would that be the final answer?

OpenStudy (amanda0618):

@rishavraj ^^

OpenStudy (amanda0618):

help me please @rishavraj

OpenStudy (amanda0618):

IS THAT THE FINAL ANSWER

OpenStudy (amanda0618):

I need someone else to help me

OpenStudy (ahsome):

x=2.57

OpenStudy (amanda0618):

THANK YOU!!

rishavraj (rishavraj):

sorry @Amanda0618 my internet connection sucks a lot......

OpenStudy (ahsome):

Its ok @rishavraj, 1+ medal

rishavraj (rishavraj):

@Ahsome thanks ...:))